# Evaluating the Use of Large Language Models as Synthetic Social Agents in Social Science Research

### v1

@@ Abstract
Large Language Models (LLMs) are being increasingly used as synthetic agents in social science, in applications ranging from augmenting survey responses to powering multi-agent simulations. Because strong prediction plus conditioning prompts, token log-probs, and repeated sampling mimic Bayesian workflows, their outputs can be misinterpreted as posterior-like evidence from a coherent model. However, prediction does not equate to probabilism, and accurate points do not imply calibrated uncertainty. This paper outlines cautions that should be taken when interpreting LLM outputs and proposes a pragmatic reframing for the social sciences in which LLMs are used as high-capacity pattern matchers for quasi-predictive interpolation under explicit scope conditions and not as substitutes for probabilistic inference. Practical guardrails such as independent draws, preregistered human baselines, reliability-aware validation, and subgroup calibration, are introduced so that researchers may engage in useful prototyping and forecasting while avoiding category errors.
@@ 1 Introduction
The advent of Large Language Models (LLMs) and their potential to act as homo silicus has sparked considerable excitement across various social science disciplines [50]. Researchers have begun leveraging these models to simulate human populations, behaviors, and social interactions [36, 39, 49, 55, 56].The growing momentum in this field suggests we may be on the cusp of a paradigm shift in social research, but it also raises urgent questions about the soundness of this approach.
A critical epistemological pitfall may lie in how we interpret LLM outputs. There must be a recognition that treating an LLM’s responses as if they were random draws from a well-defined population distribution can be deeply misleading. LLMs do generate text probabilistically by choosing each word based on learned probability distributions, but they are not probabilistic inference engines in the rigorous Bayesian sense. In orthodox predictive methods, uncertainty is quantified and propagated through models, as in Bayesian frameworks where priors are updated into posteriors given new data [48, 57]. An LLM, by contrast, does not maintain or update a posterior belief distribution when given new prompts; it has no explicit model of uncertainty or variance in the phenomena it is simulating. Its “probabilistic” output is actually the result of pre-trained static weights and sampling algorithms such as temperature or top-k/p, and not an updating mechanism that incorporates new evidence. When an LLM is instructed to act as, for example, a 30-year-old Midwestern farmer or a college-educated urban voter, it is not drawing from an actual demographic population distribution. Rather, the LLM is acting as a complex pattern-matching system that imitates what such a person might say based on traces in the LLM’s training data. If LLMs do not truly reason over distributions of people’s beliefs in the way social scientists do, then what exactly is being inferred when their outputs are treated as if they were samples from a population?
@@ 5 Making Robust Synthetic Agents
Any inference from LLM-simulated data should acknowledge distribution shift and that these are model outputs, not draws from a human population. Because models can deviate systematically from human opinions, growing a synthetic sample from itself can produce intervals that are too narrow. A better approach is to construct confidence sets that reflect mismatch and to adaptively cap the effective simulation size. [51] provide a framework that selects a data-driven size $\hat{k}$ and shows that taking k very large can hurt coverage, while the selected $\hat{k}$ maintains near-nominal coverage. This shows that alignment with human data should determine interval tightness. These types of techniques are model agnostic and use synthetic data to augment human data, not replace it. Augmentation offers a promising path forward in its own right. For example, synthetic agents have the potential to be used to pretest instruments, surface subgroup calibration issues, identify failure modes before fielding, or guide sample allocation and sensitivity analysis. In this, scarce human measurements can be collected where they are most informative. This could raise power and lower cost at the research design stage and still preserve human data as the standard for inference.

### v2

@@ Abstract
Large Language Models (LLMs) are being increasingly used as synthetic agents in social science, in applications ranging from augmenting survey responses to powering multi-agent simulations. This paper outlines cautions that should be taken when interpreting LLM outputs and proposes a pragmatic reframing for the social sciences in which LLMs are used as high-capacity pattern matchers for quasi-predictive interpolation under explicit scope conditions and not as substitutes for probabilistic inference. Practical guardrails such as independent draws, preregistered human baselines, reliability-aware validation, and subgroup calibration, are introduced so that researchers may engage in useful prototyping and forecasting while avoiding category errors.
@@ 1 Introduction
The advent of Large Language Models (LLMs) and their potential to act as homo silicus has sparked considerable excitement across various social science disciplines [50]. Researchers have begun leveraging these models to simulate human populations, behaviors, and social interactions [36, 39, 49, 55, 56].The growing momentum in this field suggests we may be on the cusp of a paradigm shift in social science methodology, but it also raises urgent questions about the soundness of this approach.
A critical epistemological danger may lie in how we interpret LLM outputs. Querying LLMs as if they operate like humans, drawing from a stable, real-world distribution, can lead to false conclusions. LLMs do generate text probabilistically by choosing each word based on learned probability distributions, but they are not probabilistic inference engines in the rigorous Bayesian sense. In orthodox predictive methods, uncertainty is quantified and propagated through models, as in Bayesian frameworks where priors are updated into posteriors given new data [48, 57]. An LLM, by contrast, does not maintain or update a posterior belief distribution when given new prompts; it has no explicit model of uncertainty or variance in the phenomena it is simulating. Its “probabilistic” output is actually the result of pre-trained static weights and sampling algorithms such as temperature or top-k/p, and not an updating mechanism that incorporates new evidence. When an LLM is instructed to act as, for example, a 30-year-old Midwestern farmer or a college-educated urban voter, it is not drawing from an actual demographic population distribution. Rather, the LLM is acting as a complex pattern-matching system that imitates what such a person might say based on traces in the LLM’s training data. If LLMs do not truly reason over distributions of people’s beliefs in the way social scientists do, then what exactly is being inferred when their outputs are treated as if they were samples from a population?
@@ 5 Making Robust Synthetic Agents
Any inference from LLM-simulated data should acknowledge distribution shift and that these are model outputs, not draws from a human population. Because models can deviate systematically from human opinions, growing a synthetic sample from itself can produce intervals that are too narrow. A better approach is to construct confidence sets that reflect mismatch and to adaptively cap the effective simulation size. [51] provide a framework that selects a data-driven size $\hat{k}$ and shows that taking k very large can hurt coverage, while the selected $\hat{k}$ maintains near-nominal coverage. This shows that alignment with human data should determine interval tightness. These types of techniques are model agnostic and use synthetic data to augment human data instead of replacing it. Augmentation offers a promising path forward in its own right. For example, synthetic agents have the potential to be used to pretest instruments, surface subgroup calibration issues, identify failure modes before fielding, or guide sample allocation and sensitivity analysis. In this, scarce human measurements can be collected where they are most informative. This could raise power and lower cost at the research design stage and still preserve human data as the standard for inference.